# C: A Coverage Path Planning Algorithm for Unknown Environments using Rapidly Covering Graphs

### v1

@@ III-B2 Progressive Sampling
The progressive sampling is the process of incremental sampling in the newly discovered area $\mathcal{A}_{d,i}$. At iteration $i\in\{0,...m-1\}$, once the robot reaches $p_{i}$ and completes the discovery step, C starts the progressive sampling step. The objective of sampling is to grow the RCG in $\mathcal{A}_{d,i}$. The RCG, in turn, generates the new waypoint $p_{i+1}$ to form the future trajectory of the robot. Note that at the final iteration $m$ there is no need for progressive sampling since the robot reaches the end point $p_{m}$ of the coverage trajectory.
a) Creation of Sampling Front: For progressive sampling, C first creates a sampling front.
@@
The sampling front $\mathcal{F}_{i}\subseteq\mathcal{A}_{d,i}$, $i\in\{0,...m-1\}$, is the obstacle-free and unsampled portion of the area $\mathcal{A}_{d,i}$ discovered in the $i^{th}$ iteration.
Figs. 5 and 6(a) show examples of the sampling front $\mathcal{F}_{i}$ generated by the robot during navigation from $p_{i-1}$ to $p_{i}$. Note that it is possible that some portion of $\mathcal{A}_{d,i}$ has been already discovered before and sampled in a previous iteration. Thus, $\mathcal{F}_{i}$ includes only the newly discovered and unsampled area.
b) Sampling Strategy: The sampling strategy is described next. Let $\mathcal{S}$ be the set of all feasible samples that lie in the coverage space $\mathcal{A}_{f}$. Once $\mathcal{F}_{i}$ is generated, it is sampled with the frontier samples $\mathcal{S}_{i}\subseteq\mathcal{S}$ as defined below.
@@
The laps are created in the sampling front $\mathcal{F}_{i}$ for the purpose of efficient, nonuniform and sparse sampling. At iteration $i\in\{0,...m-1\}$, the laps are drawn in the sampling front $\mathcal{F}_{i}$, such that they are parallel to each other and the distance between any two adjacent laps is equal to the sampling resolution $w\in\mathbb{R}^{+}$. The laps are created along the direction of desired back and forth motion. The first lap is drawn in the sampling front $\mathcal{F}_{0}$ starting from $p_{0}$. Note that it is possible that only portions of laps are created within $\mathcal{F}_{i}$, that is their ends may not reach an obstacle or the boundary. The laps are also drawn to be consistent, which means that any two laps separated by an obstacle in between, such that one lap is above and the other below the obstacle, are aligned on the same line whenever possible. Fig. 5 shows the examples of laps on the coverage trajectory and the new ones created in $\mathcal{F}_{i}$ for sampling.
@@
III-B3 Progressive RCG Growth
The objectives of a Rapidly Covering Graph (RCG) are to track the coverage progress, generate adaptive waypoints to produce the coverage trajectory through unexplored regions, and provide escape routes from the dead-ends. First, we define an RCG as follows.
@@ III-B3 Progressive RCG Growth
  • =$\mathcal{E}=\left\{e(n_{j},n_{k}):j=1,\ldots|\mathcal{N}|,k=1,\ldots|\mathcal{N}|\right\}$ is the edge set, such that each edge connects two nodes and represents a collision-free traversal path for the robot.
Definition III.6 (Neighbors).
A node $n^{\prime}\in\mathcal{N}$ is said to be a neighbor of a node $n\in\mathcal{N}$ and vice versa if there exists an edge $e(n,n^{\prime})\in\mathcal{E}$ connecting $n^{\prime}$ and $n$.
@@
The RCG is expanded as follows. First, all the frontier samples $\mathcal{S}_{i}$ in $\mathcal{F}_{i}$ are assigned to be the new nodes $\mathcal{N}_{new}$ of RCG, where $|\mathcal{N}_{new}|=|\mathcal{S}_{i}|$. The new nodes are added to the existing node set $\mathcal{N}_{i-1}$ to get $\mathcal{N}^{\prime}_{i-1}=\mathcal{N}_{i-1}\cup\mathcal{N}_{new}$. Then, the new edges $\mathcal{E}_{new}$ are created by making valid connections between new nodes and to the nodes of existing RCG. To add the edges all laps in the coverage space $\mathcal{A}_{f}$ are considered. Note that a lap in $\mathcal{A}_{f}$ may contain both new and old nodes of the existing RCG. Thus, for each new node $n\in\mathcal{N}_{new}$ on each lap, new edges are drawn to connect it with
@@
Note that it is possible that a node has more than two adjacent laps on one side separated by an obstacle. Each edge is checked for feasibility before connecting, i.e., if it lies in $\mathcal{A}_{f}$. In this manner, all possible edges are drawn from each new node that connect it to: 1) the nodes above and below on the same lap, and 2) the nodes on the left and right adjacent lap(s). The new edges are added to the existing edge set $\mathcal{E}_{i-1}$ to get $\mathcal{E}^{\prime}_{i-1}=\mathcal{E}_{i-1}\cup\mathcal{E}_{new}$. In this fashion, by adding new nodes and edges, the RCG is expanded at iteration $i$ into the sampling front $\mathcal{F}_{i}$ to get $\mathcal{G}^{\prime}_{i-1}=\{\mathcal{N}^{\prime}_{i-1},\mathcal{E}^{\prime}_{i-1}\}$.
@@
Figure 7: Flowchart to determine essential and inessential nodes.
@@
P1 and P2 directly result from the RCG expansion strategy, where the edges are only drawn from a node to connect the adjacent nodes on the same and side laps. P3 arises from the RCG expansion and pruning strategies, which build and maintain connections between all adjacent nodes on the same lap as well as connections to the ends nodes of adjacent laps, thus making every node reachable from every other node. P4 emerges from the fact that the nodes of RCG are formed from the frontier samples in the sampling front. After each pruning step only essential nodes and edges are left, and it is not possible to add any more frontier samples, thus making the RCG a minimum sufficient graph consisting of essential nodes and edges. Eventually the progressive sampling reaches the free space boundary by exploration until no more sampling front is found and no more frontier samples are created. Therefore, a fully formed RCG is a minimum sufficient graph for complete coverage. P5 appears because the RCG is created by using the frontier samples followed by the pruning process. P6 results from the Euler’s Formula [50], which states that for a simple, connected, and planar graph $|\mathcal{E}|\leq 3|\mathcal{N}|-6$.
@@ III-B4 Coverage Trajectory Generation
At each iteration $i\in\{0,...m-1\}$, once the robot reaches the waypoint $p_{i}$, generates the sampling front $\mathcal{F}_{i}$, and updates the RCG to $\mathcal{G}_{i}=\left(\mathcal{N}_{i},\mathcal{E}_{i}\right)$, the next step is to generate the next waypoint $p_{i+1}\in\mathcal{A}_{f}$. In order to generate $p_{i+1}$, one of the nodes $\hat{n}_{i+1}\in\mathcal{N}_{i}$ is selected as an intermediate goal node and its position is set as the next waypoint, such that $p_{i+1}=pos({\hat{n}_{i+1}})$, where $pos$ denotes the position of a node. The node selection procedure is described later. First, the node sequence is defined below.
@@
1 Function SelectGoalNode
2 Inputs to iteration $i$: current RCG $\mathcal{G}_{i}$; current node $\hat{n}_{i}$; $\mathcal{N}^{retreat}$.
3 if  $q(\hat{n}^{L}_{i})=Op$ then
4       $\hat{n}_{i+1}=\hat{n}^{L}_{i}$; // if left node Op, go left
@@ 7       $\hat{n}_{i+1}=\hat{n}^{U}_{i}$; // elseif up node Op, go up
9       $\hat{n}_{i+1}=\hat{n}^{D}_{i}$; // elseif down node Op, go down
@@
11       $\hat{n}_{i+1}=\hat{n}^{R}_{i}$; // elseif right node Op, go right
12 else              // else there exists a dead-end at $\hat{n}_{i}$
@@ 17            Coverage is complete;
1 Function UpdateState
2 Inputs to iteration $i$: current node $\hat{n}_{i}$; goal node $\hat{n}_{i+1}$; current node state $q(\hat{n}_{i})=Op$; $\mathcal{N}^{Op}_{i}$; $\mathcal{N}^{Cl}_{i}$.
@@ 12                   Create a link node below $\hat{n}_{i}$ at a distance $w$;
Definition III.11 (State Encoding).
At iteration $i\in\{0,...m-1\}$, once the robot reaches $\hat{n}_{i}$ (i.e., waypoint $p_{i}$) and the RCG is updated to $\mathcal{G}_{i}$, the states of all new nodes in $\mathcal{G}_{i}$ are assigned to be $Op$. Let $\mathcal{N}^{Op}_{i}\subseteq\mathcal{N}_{i}$ and $\mathcal{N}^{Cl}_{i}\subseteq\mathcal{N}_{i}$ be the sets of Open and Closed nodes of $\mathcal{G}_{i}$, respectively, such that $\mathcal{N}^{Op}_{i}\bigcup\mathcal{N}^{Cl}_{i}$$=$$\mathcal{N}_{i}$. Once the node states are updated, the goal node $\hat{n}_{i+1}$ is selected as follows.
a) Goal Node Selection Strategy: Let $\hat{n}_{i}^{L}$, $\hat{n}_{i}^{U}$, $\hat{n}_{i}^{D}$ and $\hat{n}_{i}^{R}$ be the neighbors of the current node $\hat{n}_{i}$ on the left lap, the current lap upwards, the current lap downwards, and the right lap, respectively. Note: these directions are defined with respect to a fixed coordinate frame whose vertical axis is parallel to the laps. To obtain $\hat{n}_{i+1}$ (Alg. 1 lines 3-10) an Open node is searched from the neighbors of $\hat{n}_{i}$ and selected based on the priority order as follows: $\hat{n}_{i}^{L}$rightarrow $\hat{n}_{i}^{U}$rightarrow $\hat{n}_{i}^{D}$rightarrow $\hat{n}_{i}^{R}$. In case there are more than one Open neighbors of $\hat{n}_{i}$ on the left or the right adjacent lap, then a random pick is done.
This simple node selection strategy generates the back and forth coverage pattern as follows. At any node, the robot first goes to the leftmost lap where an open node neighbor is available, then it moves on that lap until it finds another open node neighbor on the left or it reaches the end of the lap. If none of the above condition is true then it switches to the right lap and continues the coverage. In this manner, the robot performs the back and forth coverage lap by lap while shifting laps from left to right as they are covered.
Remark III.1.
b) State Update Strategy: As soon as $\hat{n}_{i+1}$ is selected, the state of $\hat{n}_{i}$ is updated as $q(\hat{n}_{i})=Cl$ (Alg. 2 lines 3-5). Only in case when both $\hat{n}^{U}_{i}\in\mathcal{N}^{Op}_{i}$ and $\hat{n}^{D}_{i}\in\mathcal{N}^{Op}_{i}$, the state of $\hat{n}_{i}$ is kept as $q(\hat{n}_{i})=Op$. This is to avoid distortion of the back and forth trajectory by marking a node as Closed in the middle of Open nodes on a lap.
If the goal is selected on the left lap, i.e., $\hat{n}_{i+1}=\hat{n}^{L}_{i}$, and $q(\hat{n}_{i})=Cl$, then it is further checked if either $\hat{n}^{U}_{i}\in\mathcal{N}^{Op}_{i}$ or $\hat{n}^{D}_{i}\in\mathcal{N}^{Op}_{i}$. If this is true then it means that the lap between $\hat{n}_{i}$ and the Open neighboring node above or below, if at a distance greater than $w$, is still uncovered. Since $q(\hat{n}_{i})=Cl$, the robot would never travel on this lap. Thus, to cover this lap in the future, a new node, called link node, is created above or below $\hat{n}_{i}$ at a distance of $w$ (Alg. 2 lines 6-13). The link nodes are pruned after they are covered and Closed.
@@
To escape from a dead-end and resume the back-and-forth coverage, an Open node needs to be found (Alg. 1 line 13). To select an open node 1) searching the entire graph is computationally inefficient and 2) randomly picking a node in uncovered region could distort the coverage trajectory. As such, C uses the concept of retreat nodes as defined below.
@@
The set $\mathcal{N}_{retreat}\subset\mathcal{N}^{Op}_{i}$ of retreat nodes is constantly updated during robot navigation. As shown in Fig. 8(a), this is done by adding any Open node to $\mathcal{N}_{retreat}$ that is within a distance of $\sqrt{2}w$ from the robot and removing any Closed node. Note that there are always some retreat nodes in the list until coverage is complete because the area is connected. If the robot reaches a dead end at $\hat{n}_{i}$, then the closest retreat node is chosen as the next goal node $\hat{n}_{i+1}$. Subsequently, the robot moves to $\hat{n}_{i+1}$ on the shortest path obtained by the A algorithm [39] and resumes the back-and-forth coverage of the remaining area. Fig. 8(b) shows an example of escaping from a dead-end to the nearest retreat node.
d) Coverage Holes Prevention Strategy: During navigation it is possible that the robot observes uncovered secluded regions that are surrounded by obstacles and covered regions. If bypassed, these regions form coverage holes, as shown in Fig. 9(a), which then require long return trajectories to cover them later. This paper presents a proactive coverage hole prevention strategy that consists of: 1) detection of potential coverage holes and 2) coverage of such coverage holes by adaptation to TSP-based locally optimal coverage trajectories.
@@
$\bullet$ Detection of Potential Coverage Holes: When the robot reaches node $\hat{n}_{i}$, it finds the next goal node $\hat{n}_{i+1}$. At this stage, it detects if there are any potential coverage holes near $\hat{n}_{i}$ that are surrounded by obstacles and covered regions. Let us index a coverage hole with $\alpha\in\mathbb{N}^{+}$ and denote the set of all nodes within this coverage hole with $\mathcal{N}^{\alpha}_{i}$.
@@
The process of finding all such coverage holes near $\hat{n}_{i}$ is as follows. First, an Open node is picked up in the neighborhood of $\hat{n}_{i}$ and assigned an index $\alpha=1$. Then, starting at this node, the neighboring Open nodes are recursively searched and labeled with index $\alpha=1$. This is done using the floodfill function. The search continues to expand in all directions until reaching Closed nodes or the goal node $\hat{n}_{i+1}$. Once the search is complete then it is further checked if any Open node with index $\alpha=1$ is adjacent to the unknown area. If true, then the set of nodes with index $\alpha=1$ is disqualified to form a coverage hole. If not, then it forms a coverage hole. Fig. 9(b) shows an example where a coverage hole is detected near $\hat{n}_{i}$. The above search is repeated for any other unlabeled Open neighbor of $\hat{n}_{i}$ and labeling it and its Open neighbors with index $\alpha=2$ until another coverage hole is detected and so on. Note that it is possible to have more than one coverage hole adjacent to $\hat{n}_{i}$. The search ends when all coverage holes near $\hat{n}_{i}$ are detected. Then $\mathcal{N}^{CH}_{i}=\bigcup_{\alpha}\mathcal{N}^{\alpha}_{i}$ denotes the set of all coverage holes around $\hat{n}_{i}$.

### v2

@@ III-B2 Progressive Sampling
The progressive sampling is the process of incremental sampling in the newly discovered area $\mathcal{A}^{d}_{i}$. At iteration $i\in\{0,...m-1\}$, once the robot reaches $p_{i}$ and completes the discovery step, C starts the progressive sampling step. The objective of sampling is to grow the RCG in $\mathcal{A}^{d}_{i}$. The RCG, in turn, generates the new waypoint $p_{i+1}$ to form the future trajectory of the robot. Note that at the final iteration $m$ there is no need for progressive sampling since the robot reaches the end point $p_{m}$ of the coverage trajectory.
a) Creation of the Sampling Front: For progressive sampling, C first creates a sampling front.
@@
The sampling front $\mathcal{F}_{i}\subseteq\mathcal{A}^{d}_{i}$, $i\in\{0,...m-1\}$, is the obstacle-free and unsampled portion of the area $\mathcal{A}^{d}_{i}$ discovered in the $i^{th}$ iteration.
Figs. 5 and 6(a) show examples of the sampling front $\mathcal{F}_{i}$ generated by the robot during navigation from $p_{i-1}$ to $p_{i}$. Note that it is possible that some portion of $\mathcal{A}^{d}_{i}$ has been already discovered before and sampled in a previous iteration. Thus, $\mathcal{F}_{i}$ includes only the newly discovered and unsampled area.
b) Sampling Strategy: The sampling strategy is described next. Let $\mathcal{S}$ be the set of all feasible samples that lie in the coverage space $\mathcal{A}^{f}$. Once $\mathcal{F}_{i}$ is generated, it is sampled with the frontier samples $\mathcal{S}_{i}\subseteq\mathcal{S}$ as defined below.
@@
The laps are created in the sampling front $\mathcal{F}_{i}$ for the purpose of efficient, nonuniform and sparse sampling. At iteration $i\in\{0,...m-1\}$, the laps are drawn in the sampling front $\mathcal{F}_{i}$, such that they are parallel to each other and the distance between any two adjacent laps is equal to the sampling resolution $w\in\mathbb{R}^{+}$. The laps are created along the direction of desired back-and-forth motion. The first lap is drawn in the sampling front $\mathcal{F}_{0}$ starting from $p_{0}$. Note that it is possible that only portions of laps are created within $\mathcal{F}_{i}$, that is their ends may not reach an obstacle or the boundary. The laps are also drawn to be consistent, which means that any two laps separated by an obstacle in between, such that one lap is above and the other below the obstacle, are aligned on the same line whenever possible. Fig. 5 shows the examples of laps on the coverage trajectory and the new ones created in $\mathcal{F}_{i}$ for sampling.
@@
III-B3 Progressive RCG Construction
The objectives of a RCG are to track the coverage progress, generate adaptive waypoints to produce the coverage trajectory through unexplored regions, and provide escape routes from the dead-ends. First, we define an RCG as follows.
@@ III-B3 Progressive RCG Construction
  • $\mathcal{E}=\left\{e(n_{j},n_{k}):j=1,\ldots|\mathcal{N}|,k=1,\ldots|\mathcal{N}|\right\}$ is the edge set, such that each edge connects two nodes and represents a collision-free traversal path for the robot.
Definition III.6 (Neighbor).
A node $n^{\prime}\in\mathcal{N}$ is said to be a neighbor of a node $n\in\mathcal{N}$ on RCG and vice versa if there exists an edge $e(n,n^{\prime})\in\mathcal{E}$ connecting $n$ and $n^{\prime}$.
@@
The RCG is expanded as follows. First, all the frontier samples $\mathcal{S}_{i}$ in $\mathcal{F}_{i}$ are assigned to be the new nodes $\mathcal{N}_{new}$ of RCG, where $|\mathcal{N}_{new}|=|\mathcal{S}_{i}|$. The new nodes are added to the existing node set $\mathcal{N}_{i-1}$ to get $\mathcal{N}^{\prime}_{i-1}=\mathcal{N}_{i-1}\cup\mathcal{N}_{new}$. Then, the new edges $\mathcal{E}_{new}$ are created by making valid connections between new nodes and to the nodes of existing RCG. To add the edges all laps in the coverage space $\mathcal{A}^{f}$ are considered. Note that a lap in $\mathcal{A}^{f}$ may contain both new and old nodes of the existing RCG. Thus, for each new node $n\in\mathcal{N}_{new}$ on each lap, new edges are drawn to connect it with
@@
Note that it is possible that a node has more than two adjacent laps on one side separated by an obstacle. Each edge is checked for feasibility before connecting, i.e., if it lies in $\mathcal{A}^{f}$. In this manner, all possible edges are drawn from each new node that connect it to: 1) the nodes above and below on the same lap, and 2) the nodes on the left and right adjacent lap(s). The new edges are added to the existing edge set $\mathcal{E}_{i-1}$ to get $\mathcal{E}^{\prime}_{i-1}=\mathcal{E}_{i-1}\cup\mathcal{E}_{new}$. In this fashion, by adding new nodes and edges, the RCG is expanded at iteration $i$ into the sampling front $\mathcal{F}_{i}$ to get $\mathcal{G}^{\prime}_{i-1}=\{\mathcal{N}^{\prime}_{i-1},\mathcal{E}^{\prime}_{i-1}\}$.
@@
Figure 7: Flowchart to determine the essential and inessential nodes.
@@
P1 and P2 directly result from the RCG expansion strategy, where the edges are only drawn from a node to connect the adjacent nodes on the same and side laps. P3 arises from the RCG expansion and pruning strategies, which build and maintain connections between all adjacent nodes on the same lap as well as connections to the ends nodes of adjacent laps, thus making every node reachable from every other node. P4 emerges from the fact that the nodes of RCG are formed from the frontier samples in the sampling front. After each pruning step only essential nodes and edges are left, and it is not possible to add any more frontier samples, thus making the RCG a minimum sufficient graph consisting of essential nodes and edges. Eventually the progressive sampling reaches the free space boundary by exploration until no more sampling front is found and no more frontier samples are created. Therefore, a fully formed RCG is a minimum sufficient graph for complete coverage. P5 appears because the RCG is created by using the frontier samples followed by the pruning process. P6 results from the Euler’s Formula [west1996introduction], which states that for a simple, connected, and planar graph $|\mathcal{E}|\leq 3|\mathcal{N}|-6$.
@@ III-B4 Coverage Trajectory Generation
At each iteration $i\in\{0,...m-1\}$, once the robot reaches the waypoint $p_{i}$, generates the sampling front $\mathcal{F}_{i}$, and updates the RCG to $\mathcal{G}_{i}=\left(\mathcal{N}_{i},\mathcal{E}_{i}\right)$, the next step is to generate the next waypoint $p_{i+1}\in\mathcal{A}^{f}$. In order to generate $p_{i+1}$, one of the nodes $\hat{n}_{i+1}\in\mathcal{N}_{i}$ is selected as an intermediate goal node and its position is set as the next waypoint, such that $p_{i+1}=pos({\hat{n}_{i+1}})$, where pospos denotes the position of a node. The node selection procedure is described later. First, the node sequence is defined below.
@@
1Function SelectGoalNode
Input: Iteration $i$: current RCG $\mathcal{G}_{i}$; current node $\hat{n}_{i}$; $\mathcal{N}^{retreat}$.
2
3if $q(\hat{n}^{L}_{i})=Op$ then
4  $\hat{n}_{i+1}=\hat{n}^{L}_{i}$; // if left node is open, go left
@@ 7  $\hat{n}_{i+1}=\hat{n}^{U}_{i}$; // if node above is open, go up
10  $\hat{n}_{i+1}=\hat{n}^{D}_{i}$; // if node down is open, go down
@@
12  $\hat{n}_{i+1}=\hat{n}^{R}_{i}$; // if right node is open, go right
13 else       // there exists a dead-end at $\hat{n}_{i}$
@@ 18    Coverage is complete;
Definition III.11 (State Encoding).
At iteration $i\in\{0,...m-1\}$, once the robot reaches $\hat{n}_{i}$ (i.e., waypoint $p_{i}$) and the RCG is updated to $\mathcal{G}_{i}$, the states of all new nodes in $\mathcal{G}_{i}$ are assigned to be $Op$. Let $\mathcal{N}^{Op}_{i}\subseteq\mathcal{N}_{i}$ and $\mathcal{N}^{Cl}_{i}\subseteq\mathcal{N}_{i}$ be the sets of Open and Closed nodes of $\mathcal{G}_{i}$, respectively, such that $\mathcal{N}^{Op}_{i}\bigcup\mathcal{N}^{Cl}_{i}$$\mathcal{N}_{i}$. Once the node states are updated, the goal node $\hat{n}_{i+1}$ is selected as follows.
a) Goal Node Selection Strategy: Let $\hat{n}_{i}^{L}$, $\hat{n}_{i}^{U}$, $\hat{n}_{i}^{D}$ and $\hat{n}_{i}^{R}$ be the neighbors of the current node $\hat{n}_{i}$ on the left lap, the current lap upwards, the current lap downwards, and the right lap, respectively. Note: these directions are defined with respect to a fixed coordinate frame whose vertical axis is parallel to the laps. To obtain $\hat{n}_{i+1}$ (Alg. 1 lines 2-9) an Open node is searched from the neighbors of $\hat{n}_{i}$ and selected based on the priority order as follows: $\hat{n}_{i}^{L}$ $\rightarrow$ $\hat{n}_{i}^{U}$ $\rightarrow$ $\hat{n}_{i}^{D}$ $\rightarrow$ $\hat{n}_{i}^{R}$. In case there are more than one Open neighbors of $\hat{n}_{i}$ on the left or the right adjacent lap, then a random pick is done.
This simple node selection strategy generates the back-and-forth coverage pattern as follows. At any node, the robot first goes to the leftmost lap where an open node neighbor is available, then it moves on that lap until it finds another open node neighbor on the left or it reaches the end of the lap. If none of the above condition is true then it switches to the right lap and continues the coverage. In this manner, the robot performs the back-and-forth coverage lap by lap while shifting laps from left to right as they are covered.
Remark III.1.
1Function UpdateState
Input: Iteration $i$: current node $\hat{n}_{i}$; goal node $\hat{n}_{i+1}$; current node state $q(\hat{n}_{i})=Op$; $\mathcal{N}^{Op}_{i}$; $\mathcal{N}^{Cl}_{i}$.
@@ 11        Create a link node below $\hat{n}_{i}$ at a distance $w$;
b) State Update Strategy: As soon as $\hat{n}_{i+1}$ is selected, the state of $\hat{n}_{i}$ is updated as $q(\hat{n}_{i})=Cl$ (Alg. 2 lines 2-4). Only in case when both $\hat{n}^{U}_{i}\in\mathcal{N}^{Op}_{i}$ and $\hat{n}^{D}_{i}\in\mathcal{N}^{Op}_{i}$, the state of $\hat{n}_{i}$ is kept as $q(\hat{n}_{i})=Op$. This is to avoid distortion of the back-and-forth trajectory by marking a node as Closed in the middle of Open nodes on a lap.
If the goal is selected on the left lap, i.e., $\hat{n}_{i+1}=\hat{n}^{L}_{i}$, and $q(\hat{n}_{i})=Cl$, then it is further checked if either $\hat{n}^{U}_{i}\in\mathcal{N}^{Op}_{i}$ or $\hat{n}^{D}_{i}\in\mathcal{N}^{Op}_{i}$. If this is true then it means that the lap between $\hat{n}_{i}$ and the Open neighboring node above or below, if at a distance greater than $w$, is still uncovered. Since $q(\hat{n}_{i})=Cl$, the robot would never travel on this lap. Thus, to cover this lap in the future, a new node, called link node, is created above or below $\hat{n}_{i}$ at a distance of $w$ (Alg. 2 lines 5-12). The link nodes are pruned after they are covered and Closed.
@@
To escape from a dead-end and resume the back-and-forth coverage, an Open node needs to be found (Alg. 1 line 12). To select an open node 1) searching the entire graph is computationally inefficient and 2) randomly picking a node in uncovered region could distort the coverage trajectory. As such, C uses the concept of retreat nodes as defined below.
@@
The set $\mathcal{N}_{retreat}\subset\mathcal{N}^{Op}_{i}$ of retreat nodes is constantly updated during robot navigation. As shown in Fig. 8(a), this is done by adding any Open node to $\mathcal{N}_{retreat}$ that is within a distance of $\sqrt{2}w$ from the robot and removing any Closed node. Note that there are always some retreat nodes in the list until coverage is complete because the area is connected. If the robot reaches a dead-end at $\hat{n}_{i}$, then the closest retreat node is chosen as the next goal node $\hat{n}_{i+1}$. Subsequently, the robot moves to $\hat{n}_{i+1}$ on the shortest path obtained by the A algorithm [hart1968formal] and resumes the back-and-forth coverage of the remaining area. Fig. 8(b) shows an example of escaping from a dead-end to the nearest retreat node.
d) Coverage Holes Prevention Strategy: During navigation it is possible that the robot observes isolated uncovered regions that are surrounded by obstacles and covered regions. If bypassed, these regions form coverage holes, as shown in Fig. 9(a), which then require long return trajectories to cover them later. This paper presents a proactive coverage hole prevention strategy that consists of: 1) detection of potential coverage holes and 2) coverage of such coverage holes by adaptation to TSP-based locally optimal coverage trajectories.
Definition III.14 (Coverage Hole).
Let $\mathcal{A}^{CH}\subseteq$ $\mathcal{A}^{uc}$ be a fully discovered but uncovered region with its boundary denoted as $\partial\mathcal{A}^{CH}$. Then, $\mathcal{A}^{CH}$ forms a coverage hole if $\forall x\in$$\partial\mathcal{A}^{CH}$, $\mathcal{B}(x,w)\setminus\mathcal{A}^{CH}\subseteq\mathcal{A}^{o}\bigcup\mathcal{A}^{c}$, i.e., it is surrounded by obstacles and covered regions.
@@
$\bullet$ Detection of Potential Coverage Holes: In order to detect coverage holes, we use RCG. When the robot reaches node $\hat{n}_{i}$, it finds the next goal node $\hat{n}_{i+1}$. At this stage, it detects if there are any potential coverage holes near $\hat{n}_{i}$ that are surrounded by obstacles and covered regions. Let us index a coverage hole around $\hat{n}_{i}$ with $\alpha\in\mathbb{N}^{+}$ and denote the set of all nodes within this coverage hole with $\mathcal{N}_{i,\alpha}^{CH}$.
@@
The process of finding all such coverage holes near $\hat{n}_{i}$ is as follows. First, an Open node is picked up in the neighborhood of $\hat{n}_{i}$ and assigned an index $\alpha=1$. Then, starting at this node, the neighboring Open nodes are recursively searched and labeled with index $\alpha=1$. This is done using the floodfill function. The search continues to expand in all directions until reaching Closed nodes or the goal node $\hat{n}_{i+1}$. Once the search is complete then it is further checked if any Open node with index $\alpha=1$ is adjacent to the unknown area. If true, then the set of nodes with index $\alpha=1$ is disqualified to form a coverage hole. If not, then it forms a coverage hole. Fig. 9(b) shows an example where a coverage hole is detected near $\hat{n}_{i}$. The above search is repeated for any other unlabeled Open neighbor of $\hat{n}_{i}$ and labeling it and its Open neighbors with index $\alpha=2$ until another coverage hole is detected and so on. Note that it is possible to have more than one coverage hole adjacent to $\hat{n}_{i}$. The search ends when all coverage holes near $\hat{n}_{i}$ are detected. Then $\mathcal{N}^{CH}_{i}=\bigcup_{\alpha}\mathcal{N}_{i,\alpha}^{CH}$ denotes the set of all coverage holes around $\hat{n}_{i}$.